The perimeter of a rectangular concrete slab is 92 feet and its area is 465 square feet. What is the length of the longer side of the slab?
Answer by friesr(113)

Perimeter is 2L + 2W
--
Area = LW
--
2l + 2w = 92
lw=465
w = 465/l
sub 465/L
2L + 2(465/l) = 92
Give it a go from there...
Once you get L, you can bring it back in for w.
Goodluck.